# Exploring a New Design Paradigm for Omnidirectional MAVs for Minimal Actuation and Internal Force Elimination: Theoretical Framework and Control

### v2

@@ V Simulation Tests
This section reports high-fidelity Gazebo simulations, implemented via the GenoM3 framework, validating the proposed controller beyond the nominal model and under more realistic conditions such as presence of parametric uncertainty. For both designs PJD1 and PJD2, the controller (10), (13) with the dynamic extension (9), shown in Fig. 3, achieves independent stabilization of the MAV attitude and position. Two indicative desired base poses are considered: an attitude command of a $60^{\circ}$ step in roll, pitch and yaw at fixed position, and a $1$ [m] position step along the inertial $(x,y,z)$ with constant orientation. The resulting trajectories are shown in Fig. 4. Simulation videos and plots for several different maneuvers are available at https://youtu.be/SWIt0CSwNUo. As noted from these simulations, the vehicle can achieve a wide range of maneuvers while generating only positive thrust, potentially reducing the complexity associated with the use of bidirectional propellers. The state and input evolutions exhibit some oscillations due to the imperfections introduced by large parametric uncertainty and actuator noise from the emulated hardware.
[…]
@@
To illustrate the stability of the zero dynamics (72) [29], as proven in Sec. -E, we simulated its evolution under nominal conditions with initial values $C^{0}_{1}(0)=I_{4\times 4},\,{\color[rgb]{0,0,0}{q_{p}}}(0)=-30^{\circ}\times\boldsymbol{1}_{6\times 1}\,$, $\dot{{\color[rgb]{0,0,0}{q_{p}}}}(0)=0.1\times\boldsymbol{1}_{6\times 1}$ rad/s. As shown in Fig. 5, the $q_{p}$ and $\dot{q}_{p}$ converge to zero, confirming convergence to the equilibrium $q_{pe}$ corresponding to that base pose $C^{0}_{1e}(0)$.
@@ VI Conclusion
Figure 7: The vehicle in Fig. 1. performing 260 degrees roll.
This work proposes a novel omnidirectional MAV concept enabling full 6-DoF motion with a minimum number of actuators and no internal forces at equilibrium. The design combines a rigid base evolving on $SE(3)$, thus avoiding representation singularities, with 4 propeller-carrying links connected through universal joints, 3 of which are passive whereas 1 is actively driven by 2 servomotors. We proved that the proposed MAV satisfies the Input-Dimension, Equilibria, Internal-Force, and Stabilizability properties. Without overactuation, an equilibrium exists for any base pose, and when torsion springs are employed, the total thrust exactly balances the vehicle weight. Even without springs, internal forces can be made arbitrarily small through changing the links masses and lengths. Any base pose can be rendered locally asymptotically stable via a dynamic FBL and backstepping from base pose error on $SE(3)$. High-fidelity simulations demonstrate decoupled position and attitude control under realistic conditions.
A possible mechanical implementation is illustrated in Fig. 7. This design scales favorably, as hovering loads are distributed almost evenly among propellers, avoiding excessive stress on the actively tilting unit. Scaling challenges are therefore typical to those of multirotors, such as propeller and battery sizing, rather than limitations specific to the proposed MAV.
Future work will focus on fully geometric control formulations on $SE(3)\times\mathbb{T}^{8}$, exploiting the passive-joint dynamics to construct a passivity-based controller, investigating a more generic propeller placements, extending the scheme to handle trajectory tracking, and carrying out experimental validation.

### v3

@@ V Simulation Tests
This section reports high-fidelity Gazebo simulations, implemented via the GenoM3 framework, validating the proposed controller beyond the nominal model and under more realistic conditions such as presence of parametric uncertainty. For both designs PJD1 and PJD2, the controller (10), (13) with the dynamic extension (9), shown in Fig. 3, achieves independent stabilization of the MAV attitude and position. Two indicative desired base poses are considered: an attitude command of a $60^{\circ}$ step in roll, pitch and yaw at fixed position, and a $1$ [m] position step along the inertial $(x,y,z)$ with constant orientation. The resulting trajectories are shown in Fig. 4. Simulation videos and plots for several different maneuvers are available at https://www.youtube.com/playlist?list=PL4N8pJgvqASQX6AWEpg3NCZ6QdGBPfbXq. Most notably, a simulation with $270$-degree roll turn is included (see also Fig. 7and 8 [29]) which demonstrates full omnidirectional capabilities. As it can be noted from these simulations, the vehicle can achieve a wide range of maneuvers while generating only positive thrust, potentially reducing the complexity associated with the use of bidirectional propellers. The state and input evolutions exhibit the natural real-world-like oscillations due to the imperfections introduced by parametric uncertainty and actuator noise from the emulated hardware.
@@
[…]
To illustrate the stability of the zero dynamics (72[29], as proven in Sec. -E, we simulated its evolution under nominal conditions with initial values $C^{0}_{1}(0)=I_{4\times 4},\,{\color[rgb]{0,0,0}{q_{p}}}(0)=-30^{\circ}\times\boldsymbol{1}_{6\times 1}\,$, $\dot{{\color[rgb]{0,0,0}{q_{p}}}}(0)=0.1\times\boldsymbol{1}_{6\times 1}$ rad/s. As shown in Fig. 5, the $q_{p}$ and $\dot{q}_{p}$ converge to zero, confirming convergence to the equilibrium $q_{pe}$ corresponding to that base pose $C^{0}_{1e}(0)$.
@@ VI Conclusion
Figure 7: Top: the PenduMAV, simulated with an independent and uncertain/noisy model in Gazebo with the DART physics engine, performing a 270-degree roll maneuver. Video with different maneuvers available at https://www.youtube.com/playlist?list=PL4N8pJgvqASQX6AWEpg3NCZ6QdGBPfbXq.
Bottom: Position and attitude time evolutions of the two phases.
This work proposes the PenduMAV a novel omnidirectional MAV concept enabling full 6-DoF motion with a minimum number of actuators and no internal forces at equilibrium. The design combines a rigid base evolving on $SE(3)$, thus avoiding representation singularities, with 4 propeller-carrying links connected through universal joints, 3 of which are passive whereas 1 is actively driven by 2 servomotors. We proved that the proposed MAV satisfies the Input-Dimension, Equilibria, Internal-Force, and Stabilizability properties. Without overactuation, an equilibrium exists for any base pose, and when torsion springs are employed, the total thrust exactly balances the vehicle weight. Even without springs, internal forces can be made arbitrarily small through changing the links masses and lengths. Any base pose can be rendered locally asymptotically stable via a dynamic FBL and backstepping from base pose error on $SE(3)$. High-fidelity simulations demonstrate decoupled position and attitude control under realistic conditions.
This design scales favorably, as hovering loads are distributed almost evenly among propellers, avoiding excessive stress on the actively tilting unit. Scaling challenges are therefore typical to those of multirotors, such as propeller and battery sizing, rather than limitations specific to the proposed MAV. Future works will focus on a large and long term project to realize, implement and test the mechanical concept.